Which label describes the following?
3x + 6 = 10
A) Linear equation B) Equation which is not linear C) Expression


Isn't it B!

Answer by tran3209(100)

3x+6=10
3x=10-6
3x=4
x=4/3
This is a linear equation of a line perpendicular to axis of x at the point (+4/3 0)
The answer is A